Vertex (h,k) of the quadratic function f(x)=a(x-h)^+k is (1,2) and the y-intercept is (0,1).
a)Find the quadratic function
b)What is the maximum value of the function?

Answer by stanbon(75887)

Vertex (h,k) of the quadratic function f(x)=a(x-h)^+k is (1,2) and the y-intercept is (0,1).
a)Find the quadratic function
Using the vertex values you get:
y = a(x-1)^2 + 2
Using the y-intercet you get:
1 = a(0-1)^2+2
1 = a+2
a = -1
---------
EQUATION:
f(x) = -(x-1)^2+2
------------------------
b)What is the maximum value of the function?
The vertex is (1,2) and the parabola opens downward,
so the maximum value of the function is 2.
==================
Cheers,
Stan H.